For 10 observations on price (X) and supply (Y) the following data were

obtained (in appropriate units):

∑X = 130, ∑Y = 220, ∑X


2 = 2288, ∑Y


2 = 5506 and ∑XY = 3467.


Obtain the line of regression of Y on X and estimate the supply when the price is 16

units, and find out the standard error of the estimate.


1
Expert's answer
2021-12-23T17:56:36-0500

The line of regression of Y on X is expressed as:

"Y=a+bX"

Where;

"b=\frac{n\Sigma XY-(\Sigma X)(\Sigma Y)}{n \Sigma X^2-(\Sigma X)^2}"


"b=\frac{(10\times3467)-(130)(220)}{10 (2288)^2-(130)^2}"

"b=1.02"


"a=\frac{\Sigma Y-b\Sigma X}{n}"


"a=\frac{220-1.02\times130}{10}"


"a=8.74"


"Y=8.74+1.02X"


If the price (X) is 16 units then supply (Y) is:


"Y=8.74+1.02\times16"


"Y=25.06"


The standard error of the estimate is:


"s=\sqrt{\frac{\Sigma Y^2-a\Sigma Y-b\Sigma XY}{n-2}}"


"s=\sqrt{\frac{5506-(8.74\times220)-(1.02\times3467)}{10-2}}"


"s=2.42"